1.      The roots of the equation x²  - 6x + I = 0 are?

2.      The equation of the line of symmetry to the curve y =3x²  - 6x + 5 = 0 is?


1
Expert's answer
2021-02-25T05:06:04-0500

1)x2 - 6x + 1 = 0

D = (-6)2 - 4·1·1 = 36 - 4 = 32

x1 = 6 - √322·1 = 3 - 2√2 ≈ 0.172

x2 = 6 + √322·1 = 3 + 2√2 ≈ 5.828


2) y =3x² - 6x + 5 = 0 

if a=3 and b=-6 let's use the next formula:

x= -b/2a=-(-6)/2*3=-(-1)=1, so

x=1 and y=0

he equation of the line of symmetry (1;0) with next form:

1x +0y-1=0